OpenStudy (anonymous):

If f(x) = 2x2 + 4, find f(3).

OpenStudy (anonymous):

Plug 3 in every where you see an "x" and evaluate.

OpenStudy (anonymous):

f(3)=2x3x2+4=12+4=16